### v2

@@ Abstract.
Spatial-temporal causal time series (STC-TS) involve region-specific temporal observations driven by causally relevant covariates and interconnected across geographic or network-based spaces. Existing methods often model spatial and temporal dynamics independently and overlook causality-driven probabilistic forecasting, limiting their predictive power. To address this, we propose STOAT (Spatial-Temporal Probabilistic Causal Inference Network), a novel framework for probabilistic forecasting in STC-TS. The proposed method extends a causal inference approach by incorporating a spatial relation matrix that encodes interregional dependencies (e.g. proximity or connectivity), enabling spatially informed causal effect estimation. The resulting latent series are processed by deep probabilistic models to estimate the parameters of the distributions, enabling calibrated uncertainty modeling. We further explore multiple output distributions (e.g., Gaussian, Student’s-$t$, Laplace) to capture region-specific variability. Experiments on COVID-19 data across six countries demonstrate that STOAT outperforms state-of-the-art probabilistic forecasting models (DeepAR, DeepVAR, Deep State Space Model, etc.) in key metrics, particularly in regions with strong spatial dependencies. By bridging causal inference and geospatial probabilistic forecasting, STOAT offers a generalizable framework for complex spatial-temporal tasks, such as epidemic management.
@@
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Proceedings of the 33rd ACM SIGSPATIAL International Conference on Advances in Geographic Information Systems; November 03–06, 2025; Minneapolis, MN, USAisbn: 978-1-4503-XXXX-X/2025/11
@@ 1. Introduction
Spatial-temporal causal time series (STC-TS) consist of regional time series that are influenced by causally relevant covariates and interconnected through spatial dependencies, as shown in Figure 1. A representative case is epidemic forecasting across countries, where region-level outcomes such as case counts or hospitalizations are influenced by external factors including government interventions, vaccination rates, mobility patterns, population density, and virus evolution (Cao and Liu, 2022; Cao, 2022; Nolan et al., 2020; Du et al., 2025). These covariates are heterogeneous, time-varying, and exhibit complex causal impacts on the temporal trajectory of the target variables. For epidemic forecasting, the target variables such as confirmed or hospitalized cases form region-specific time series, while external covariates evolve alongside, often asynchronously and irregularly. The complex causal interactions between such covariates (for instance, the joint effect of virus mutations and policy responses) remain difficult to model explicitly, particularly when considering their varying impacts across different regions (Mastakouri and Schölkopf, 2020). Furthermore, regional outcomes are not isolated phenomena. Spatial dependencies arise naturally as outbreak patterns often exhibit spillover effects due to cross-border mobility, trade connections, and latent transmission pathways (Wang et al., 2024; Salim, 2021). Understanding spatio-temporal dynamics therefore requires modeling both the causal effects of covariates within each region and the spatial influences propagating between regions. These interconnected challenges in STC-TS settings give rise to three fundamental modeling requirements: cross-region causal inference, spatial-temporal causal representation learning, and causally informed probabilistic forecasting.
@@
First Challenge: Cross-region causal inference. Factors and events across different regions (such as policy changes, mobility shifts, and vaccination progress) often have interdependent spillover effects on the outcome variables through spatial (Naumann et al., 2020). However, existing deep probabilistic forecasting models do not explicitly model how these cross-regional interactions and spatial dependencies jointly affect target dynamics across multiple regions. For example, DeepAR (Salinas et al., 2020) and Deep State Space Models (Rangapuram et al., 2018) treat covariates as auxiliary inputs without capturing spatial causal relationships between regions, lacking interpretability in understanding how interventions propagate across spatial networks. Classical Difference-in-Differences (DiD) (Ryan et al., 2019; Metcalfe et al., 2019; Rothbard et al., 2024) enables causal inference but is typically designed for simple treatment-control setups without considering spatial spillover effects or the complex interdependencies inherent in spatial networks. In contrast to these approaches, our framework explicitly models spatial causal mechanisms through a spatial lag term that quantifies how neighboring regions’ outcomes causally influence each target region, while simultaneously incorporating treatment effects within a unified causal inference structure. Bridging this gap by developing spatially-aware causal inference frameworks that can both capture cross-regional spillover effects and provide interpretable causal estimates for complex spatio-temporal time series represents a critical yet largely unaddressed challenge in the intersection of causal inference and deep learning.
@@
Second Challenge: Spatial-Temporal Causal Representation Learning. Classical approaches in epidemiology, including compartmental models (e.g., SIR (Cooper et al., 2020)) and statistical methods such as Difference-in-Differences (Ohlsson and Kendler, 2020; Tchetgen et al., 2024), Instrumental Variables (IV) (Baiocchi et al., 2014), and Propensity Score Matching (PSM) (Kane et al., 2020), lack the capacity to capture high-dimensional, non-linear, and dynamic causal dependencies. These methods typically assume static confounding structures and fail to model complex spatio-temporal interactions in a data-driven manner. Meanwhile, existing deep temporal models often treat covariates as exogenous inputs without accounting for structural confounding or causal entanglement. Spatial dependencies, if considered at all, are rarely modeled through a causal lens. Consequently, both classical and modern approaches remain limited in their ability to infer structured causal dynamics over time and across regions. To address this, we argue that accurate forecasting in spatial-temporal causal time series (STC-TS) requires latent representations that encode both temporal dependencies and spatially heterogeneous causal effects.
@@
Third Challenge: Causally Informed Probabilistic Forecasting. Causally informed probabilistic forecasting is critical in high-stakes scenarios like epidemic control, where uncertainty quantification is as important as point predictions (Xie et al., 2023). However, existing methods often rely on deterministic predictions and fail to adaptively model the distributional dynamics induced by time-varying covariates or interventions, overlooking the causal relationships between covariates and targets (Gneiting and Katzfuss, 2014). As a result, predictive intervals are miscalibrated and insensitive to policy changes, regional disparities, and intervention timing. Furthermore, these models typically do not distinguish between the different sources of uncertainty, such as aleatoric, epistemic, and structural uncertainty, which limits both interpretability and the robustness of downstream decision-making.
@@
To address the unique challenges of spatial-temporal causal time series (STC-TS) forecasting, we introduce STOAT, a unified framework that integrates causal inference, spatial-temporal representation learning, and probabilistic forecasting. STOAT is built on the premise that reliable forecasting under covariate interventions and spatial spillovers requires more than standard temporal modeling; it requires structured causal reasoning and calibrated uncertainty estimation. The key design principle behind STOAT is to first disentangle causal effects from spatio-temporal confounding through a dedicated Spatial-Temporal Causal Inference Module, which extends the classical Difference-in-Differences (DiD) (Ryan et al., 2019; Metcalfe et al., 2019; Rothbard et al., 2024) using a learnable spatial relation matrix. This module yields causally adjusted representations that reflect both treatment effects and spatial spillovers across regions. These representations are then fed into a Deep Probabilistic Forecasting Module, which captures temporal dependencies using neural sequence models and projects the latent states into structured output distributions. This enables STOAT to produce calibrated probabilistic forecasts that account for both intervention dynamics and spatial correlations. Figure 2 provides an overview of the STOAT architecture, which consists of two interconnected modules: (1) a spatial-temporal causal inference pathway that generates covariate-adjusted representations, and (2) a probabilistic forecasting pathway that learns temporal dynamics and generates distributional predictions. The primary contributions of this work are summarized as follows:
@@
Figure 1. Illustration of the Spatial-temporal causal time series (STC-TS) setting. The bottom-left pink block (causal dependencies) depicts intra-regional causal inference, where time-varying covariates (e.g., policy interventions) influence observed outcomes within each region. Black-and-white nodes represent causally relevant covariates at specific time steps. The top gray block (temporal dependencies) shows the evolution of observed variables across three regions, with colored points (red, yellow, green) denoting distinct temporal trends. The bottom-right green block (spatial dependencies) illustrates inter-regional relationships, highlighting spatially structured influences and spillover effects. Together, the figure captures the intertwined spatial, temporal, and causal structures that define the STC-TS framework.
[…]
A conceptual diagram showing three interconnected blocks representing the STC-TS framework: a pink block showing causal dependencies with nodes and arrows, a gray block displaying temporal evolution across regions with colored trend lines, and a green block illustrating spatial relationships between regions.
@@ 2. Related Work
Causal Inference in Time Series. Causal inference has been extensively studied in econometrics and policy analysis, especially through the framework of Difference-in-Differences (DiD) (Abadie, 2005; Moraffah et al., 2021; Athey and Imbens, 2017; Runge et al., 2023). DiD has been widely applied to time series data to estimate treatment effects over time by comparing treated and control groups. Recent extensions include generalized DiD methods for staggered treatment adoption (Sun and Abraham, 2021) and synthetic control-based approaches (Abadie et al., 2010). However, these methods are typically restricted to one-dimensional or pairwise group comparisons, and do not account for spatial interactions across regions. They also assume static treatment effects and rarely integrate with representation learning. While some recent work incorporates DiD with neural networks (Zhou et al., 2024; Schwab et al., 2020), they lack structured spatial modeling and do not support multi-region causal adjustment. Our work addresses this gap by integrating spatial causal inference into a deep temporal forecasting framework using a structured adjacency matrix for regional spillover modeling.
@@
Deep Probabilistic Time-Series Forecasting. Several deep learning models have been proposed for probabilistic time series forecasting by combining recurrent neural networks with statistical likelihood estimation. DeepAR (Salinas et al., 2020) and MTSNet (Yang and Cao, 2023) model each time series independently under a univariate distribution, using shared RNN parameters to learn sequential dependencies. Deep State Space Models (DSSM) (Rangapuram et al., 2018) incorporate a linear state-space formulation with recurrent updates, but similarly assume independent marginal distributions for each series. DeepVAR (Salinas et al., 2019) enhances this with Gaussian copula processes to capture dependencies between target variables but does not model structured causal relationships across space or time. DeepFactor (Wang et al., 2019) decomposes global and local components using latent variables, yet still treats covariates as exogenous inputs without causal interpretation. None of these models incorporate causal inference mechanisms or address confounding in the presence of spatially distributed covariates. In contrast, STOAT explicitly adjusts for covariate-driven biases through spatial causal representations and generates calibrated predictive distributions informed by intervention dynamics.
@@
Sequence Modeling for Epidemic Forecasting. Sequence-to-sequence (seq2seq) models and encoder-decoder architectures have been widely applied in epidemic forecasting (Pathan et al., 2020; Alassafi et al., 2022; Shahid et al., 2020). These models, including GRU-based and transformer variants, are primarily trained for point prediction, optimizing loss functions such as MSE or MAE. While effective in short-term forecasting, they do not quantify predictive uncertainty or incorporate causal reasoning. Classical epidemiological models, such as SIR and SEIR variants (Cao and Liu, 2022; He et al., 2020; Carcione et al., 2020; López and Rodo, 2021), provide mechanistic insights into disease transmission, but they rely on fixed parametric assumptions and are difficult to scale or integrate with heterogeneous covariates. Moreover, these models often do not support flexible uncertainty modeling or data-driven representation learning. Our work bridges this gap by combining causal inference with probabilistic deep learning. STOAT allows for region-aware, covariate-adjusted forecasting with interpretable uncertainty, offering a unified approach to modeling spatial-temporal dynamics in complex real-world settings. Compared to existing approaches, STOAT differs in three key aspects:
@@ 3.1. Overview of STOAT
Spatial-temporal causal time series (STC-TS) consisting of target variables time-series $\mathbf{y}$ and multiple causally relevant covariates time-series $\mathbf{C}$ exhibit both inter-regional spatial dependencies and intra-regional temporal couplings, as shown in Figure 1. For the considered epidemic forecasting data, we assume $\mathbf{y}_{t}$ represents a target vector composed of the values of the multivariate time series at time step $t$, e.g., confirmed cases from $N$ regions/countries. $\mathbf{C}_{t}$ represents the causally relevant covariates corresponding to the target multivariate time series, e.g., policy stringency index (Hale et al., 2021; Gros et al., 2021), vaccination rates, ICU patient and reproduction number (Alimohamadi et al., 2020) from $N$ regions/countries, where these covariates serve as causal factors (treatments) that influence the target outcomes. STOAT characterizes the STC-TS probabilistic forecasting problem in Eq. (1),
@@ 3.2. Spatial-Temporal Causal Inference Module
The Spatial-Temporal Causal Inference Module performs causally-informed representation learning by extending the Difference-in-Differences (Moraffah et al., 2021) framework with spatial relation matrices, enabling region-aware causal adjustment representation learning.
@@ 3.2.2. Spatial Causal Inference
(3) $y_{i,t}=\rho\sum_{j=1}^{N}S_{i,j}y_{j,t}+\beta_{0}+\beta_{1}T_{i}+\beta_{2}\text{Post}_{t}+\delta(T_{i}\cdot\text{Post}_{t})+\boldsymbol{\gamma}^{\top}\mathbf{c}_{i,t}+\epsilon_{i,t}$
where $y_{i,t}$ represents the observed outcome (e.g., COVID-19 cases) for region $i$ at time $t$. The spatial lag term $\sum_{j=1}^{N}S_{i,j}y_{j,t}$ captures the weighted average of neighboring regions’ outcomes based on the spatial relation matrix, with coefficient $\rho$ controlling the strength of spatial spillover effects. The treatment indicator $T_{i}$ is a binary variable denoting whether region $i$ has adopted the intervention, while $\text{Post}_{t}$ indicates the post-treatment period. The parameter $\delta$ represents the treatment effect, capturing the causal impact of interventions on target outcomes.

### v3

@@ Abstract.
Spatial-temporal causal time series (STC-TS) involve region-specific temporal observations driven by causally relevant covariates and interconnected across geographic or network-based spaces. Existing methods often model spatial and temporal dynamics independently and overlook causality-driven probabilistic forecasting, limiting their predictive power. To address this, we propose STOAT (Spatial-Temporal Probabilistic Causal Inference Network), a novel framework for probabilistic forecasting in STC-TS. The proposed method extends a causal inference approach by incorporating a spatial relation matrix that encodes interregional dependencies (e.g. proximity or connectivity), enabling spatially informed causal effect estimation. The resulting latent series are processed by deep probabilistic models to estimate the parameters of the distributions, enabling calibrated uncertainty modeling. We further explore multiple output distributions (e.g., Gaussian, Student-t, Laplace) to capture region-specific variability. Experiments on COVID-19 data across six countries demonstrate that STOAT outperforms state-of-the-art probabilistic forecasting models (DeepAR, DeepVAR, Deep State Space Model, etc.) in key metrics, particularly in regions with strong spatial dependencies. By bridging causal inference and geospatial probabilistic forecasting, STOAT offers a generalizable framework for complex spatial-temporal tasks, such as epidemic management.
@@
copyright: acmlicensedjournalyear: 2025doi: 10.1145/3748636.3762761conference: Proceedings of the 33rd ACM SIGSPATIAL International Conference on Advances in Geographic Information Systems; November 03–06, 2025; Minneapolis, MN, USAisbn: 979-8-4007-2086-4/2025/11
@@ 1. Introduction
Spatial-temporal causal time series (STC-TS) consist of regional time series that are influenced by causally relevant covariates and interconnected through spatial dependencies, as shown in Figure 1. A representative case is epidemic forecasting across countries, where region-level outcomes such as case counts or hospitalizations are influenced by external factors including government interventions, vaccination rates, mobility patterns, population density, and virus evolution (cao2021covid; cao2022ai; nolan2020exploring; du2025blue). These covariates are heterogeneous, time-varying, and exhibit complex causal impacts on the temporal trajectory of the target variables. For epidemic forecasting, the target variables such as confirmed or hospitalized cases form region-specific time series, while external covariates evolve alongside, often asynchronously and irregularly. The complex causal interactions between such covariates (for instance, the joint effect of virus mutations and policy responses) remain difficult to model explicitly, particularly when considering their varying impacts across different regions (mastakouri2020causal). Furthermore, regional outcomes are not isolated phenomena. Spatial dependencies arise naturally as outbreak patterns often exhibit spillover effects due to cross-border mobility, trade connections, and latent transmission pathways (wang2024learning; salim2021learning). Understanding spatio-temporal dynamics therefore requires modeling both the causal effects of covariates within each region and the spatial influences propagating between regions. These interconnected challenges in STC-TS settings give rise to three fundamental modeling requirements: cross-region causal inference, spatial-temporal causal representation learning, and causally informed probabilistic forecasting.
@@
First Challenge: Cross-region causal inference. Factors and events across different regions (such as policy changes, mobility shifts, and vaccination progress) often have interdependent spillover effects on the outcome variables through spatial (naumann2020covid). However, existing deep probabilistic forecasting models do not explicitly model how these cross-regional interactions and spatial dependencies jointly affect target dynamics across multiple regions. For example, DeepAR (salinas2020deepar) and Deep State Space Models (rangapuram2018deep) treat covariates as auxiliary inputs without capturing spatial causal relationships between regions, lacking interpretability in understanding how interventions propagate across spatial networks. Classical Difference-in-Differences (DiD) (ryan2019now; metcalfe2019pay; rothbard2024tutorial) enables causal inference but is typically designed for simple treatment-control setups without considering spatial spillover effects or the complex interdependencies inherent in spatial networks. In contrast to these approaches, our framework explicitly models spatial causal mechanisms through a spatial lag term that quantifies how neighboring regions’ outcomes causally influence each target region, while simultaneously incorporating treatment effects within a unified causal inference structure. Bridging this gap by developing spatially-aware causal inference frameworks that can both capture cross-regional spillover effects and provide interpretable causal estimates for complex spatio-temporal time series represents a critical yet largely unaddressed challenge in the intersection of causal inference and deep learning.
@@
Second Challenge: Spatial-Temporal Causal Representation Learning. Classical approaches in epidemiology, including compartmental models (e.g., SIR (cooper2020sir)) and statistical methods such as Difference-in-Differences (ohlsson2020applying; tchetgen2024universal), Instrumental Variables (IV) (baiocchi2014instrumental), and Propensity Score Matching (PSM) (kane2020propensity), lack the capacity to capture high-dimensional, non-linear, and dynamic causal dependencies. These methods typically assume static confounding structures and fail to model complex spatio-temporal interactions in a data-driven manner. Meanwhile, existing deep temporal models often treat covariates as exogenous inputs without accounting for structural confounding or causal entanglement. Spatial dependencies, if considered at all, are rarely modeled through a causal lens. Consequently, both classical and modern approaches remain limited in their ability to infer structured causal dynamics over time and across regions. To address this, we argue that accurate forecasting in spatial-temporal causal time series (STC-TS) requires latent representations that encode both temporal dependencies and spatially heterogeneous causal effects.
@@
Third Challenge: Causally Informed Probabilistic Forecasting. Causally informed probabilistic forecasting is critical in high-stakes scenarios like epidemic control, where uncertainty quantification is as important as point predictions (xie2023overview). However, existing methods often rely on deterministic predictions and fail to adaptively model the distributional dynamics induced by time-varying covariates or interventions, overlooking the causal relationships between covariates and targets (gneiting2014probabilistic). As a result, predictive intervals are miscalibrated and insensitive to policy changes, regional disparities, and intervention timing. Furthermore, these models typically do not distinguish between the different sources of uncertainty, such as aleatoric, epistemic, and structural uncertainty, which limits both interpretability and the robustness of downstream decision-making.
@@
To address the unique challenges of spatial-temporal causal time series (STC-TS) forecasting, we introduce STOAT, a unified framework that integrates causal inference, spatial-temporal representation learning, and probabilistic forecasting. STOAT is built on the premise that reliable forecasting under covariate interventions and spatial spillovers requires more than standard temporal modeling; it requires structured causal reasoning and calibrated uncertainty estimation. The key design principle behind STOAT is to first disentangle causal effects from spatio-temporal confounding through a dedicated Spatial-Temporal Causal Inference Module, which extends the classical Difference-in-Differences (DiD) (ryan2019now; metcalfe2019pay; rothbard2024tutorial) using a learnable spatial relation matrix. This module yields causally adjusted representations that reflect both treatment effects and spatial spillovers across regions. These representations are then fed into a Deep Probabilistic Forecasting Module, which captures temporal dependencies using neural sequence models and projects the latent states into structured output distributions. This enables STOAT to produce calibrated probabilistic forecasts that account for both intervention dynamics and spatial correlations. Figure 2 provides an overview of the STOAT architecture, which consists of two interconnected modules: (1) a spatial-temporal causal inference pathway that generates covariate-adjusted representations, and (2) a probabilistic forecasting pathway that learns temporal dynamics and generates distributional predictions. The primary contributions of this work are summarized as follows:
@@
Figure 1. Illustration of the Spatial-temporal causal time series (STC-TS) setting. The bottom-left pink block (causal dependencies) depicts intra-regional causal inference, where time-varying covariates (e.g., policy interventions) influence observed outcomes within each region. Black-and-white nodes represent causally relevant covariates at specific time steps. The top gray block (temporal dependencies) shows the evolution of observed variables across three regions, with colored points (red, yellow, green) denoting distinct temporal trends. The bottom-right green block (spatial dependencies) illustrates inter-regional relationships, highlighting spatially structured influences and spillover effects. Together, the figure captures the intertwined spatial, temporal, and causal structures that define the STC-TS framework.:
@@ 2. Related Work
Causal Inference in Time Series. Causal inference has been extensively studied in econometrics and policy analysis, especially through the framework of Difference-in-Differences (DiD) (abadie2005semiparametric; moraffah2021causal; athey2017state; runge2023causal). DiD has been widely applied to time series data to estimate treatment effects over time by comparing treated and control groups. Recent extensions include generalized DiD methods for staggered treatment adoption (sun2021estimating) and synthetic control-based approaches (abadie2010synthetic). However, these methods are typically restricted to one-dimensional or pairwise group comparisons, and do not account for spatial interactions across regions. They also assume static treatment effects and rarely integrate with representation learning. While some recent work incorporates DiD with neural networks (zhou2024machine; schwab2020learning), they lack structured spatial modeling and do not support multi-region causal adjustment. Our work addresses this gap by integrating spatial causal inference into a deep temporal forecasting framework using a structured adjacency matrix for regional spillover modeling.
@@
Deep Probabilistic Time-Series Forecasting. Several deep learning models have been proposed for probabilistic time series forecasting by combining recurrent neural networks with statistical likelihood estimation. DeepAR (salinas2020deepar) and MTSNet (yang2023mtsnet) model each time series independently under a univariate distribution, using shared RNN parameters to learn sequential dependencies. Deep State Space Models (DSSM) (rangapuram2018deep) incorporate a linear state-space formulation with recurrent updates, but similarly assume independent marginal distributions for each series. DeepVAR (salinas2019high) enhances this with Gaussian copula processes to capture dependencies between target variables but does not model structured causal relationships across space or time. DeepFactor (wang2019deep) decomposes global and local components using latent variables, yet still treats covariates as exogenous inputs without causal interpretation. None of these models incorporate causal inference mechanisms or address confounding in the presence of spatially distributed covariates. In contrast, STOAT explicitly adjusts for covariate-driven biases through spatial causal representations and generates calibrated predictive distributions informed by intervention dynamics.
@@
Sequence Modeling for Epidemic Forecasting. Sequence-to-sequence (seq2seq) models and encoder-decoder architectures have been widely applied in epidemic forecasting (pathan2020time; alassafi2022time; shahid2020predictions). These models, including GRU-based and transformer variants, are primarily trained for point prediction, optimizing loss functions such as MSE or MAE. While effective in short-term forecasting, they do not quantify predictive uncertainty or incorporate causal reasoning. Classical epidemiological models, such as SIR and SEIR variants (cao2021covid; he2020seir; carcione2020simulation; lopez2021modified), provide mechanistic insights into disease transmission, but they rely on fixed parametric assumptions and are difficult to scale or integrate with heterogeneous covariates. Moreover, these models often do not support flexible uncertainty modeling or data-driven representation learning. Our work bridges this gap by combining causal inference with probabilistic deep learning. STOAT allows for region-aware, covariate-adjusted forecasting with interpretable uncertainty, offering a unified approach to modeling spatial-temporal dynamics in complex real-world settings. Compared to existing approaches, STOAT differs in three key aspects:
@@ 3.1. Overview of STOAT
Spatial-temporal causal time series (STC-TS) consisting of target variables time-series $\mathbf{y}$ and multiple causally relevant covariates time-series $\mathbf{C}$ exhibit both inter-regional spatial dependencies and intra-regional temporal couplings, as shown in Figure 1. For the considered epidemic forecasting data, we assume $\mathbf{y}_{t}$ represents a target vector composed of the values of the multivariate time series at time step $t$, e.g., confirmed cases from $N$ regions/countries. $\mathbf{C}_{t}$ represents the causally relevant covariates corresponding to the target multivariate time series, e.g., policy stringency index (hale2021global; gros2021new), vaccination rates, ICU patient and reproduction number (alimohamadi2020estimate) from $N$ regions/countries, where these covariates serve as causal factors (treatments) that influence the target outcomes. STOAT characterizes the STC-TS probabilistic forecasting problem in Eq. (1),
@@ 3.2. Spatial-Temporal Causal Inference Module
The Spatial-Temporal Causal Inference Module performs causally-informed representation learning by extending the Difference-in-Differences (moraffah2021causal) framework with spatial relation matrices, enabling region-aware causal adjustment representation learning.
@@ 3.2.2. Spatial Causal Inference
(3) $y_{i,t}=\rho\sum_{j=1}^{N}S_{i,j}y_{j,t-1}+\beta_{0}+\beta_{1}T_{i}+\beta_{2}\text{Post}_{t}+\delta(T_{i}\cdot\text{Post}_{t})+\boldsymbol{\gamma}^{\top}\mathbf{c}_{i,t}+\epsilon_{i,t}$
where $y_{i,t}$ represents the observed outcome (e.g., COVID-19 cases) for region $i$ at time $t$. The spatial lag term $\sum_{j=1}^{N}S_{i,j}y_{j,t-1}$ captures the weighted average of neighboring regions’ outcomes based on the spatial relation matrix, with coefficient $\rho$ controlling the strength of spatial spillover effects. The treatment indicator $T_{i}$ is a binary variable denoting whether region $i$ has adopted the intervention, while $\text{Post}_{t}$ indicates the post-treatment period. The parameter $\delta$ represents the treatment effect, capturing the causal impact of interventions on target outcomes.